When 2.00 liters of SO3 gas at STP react, how many grams of O2 gas will form?
2SO3(g) → 2SO2(g) + O2(g)
1
Expert's answer
2020-10-26T13:54:54-0400
n(SO3)= V/Vm      n(SO3)=2/22,2=0,09 mole
n(O2)=0,089/2=0,045 mole

"m=n*M"

m(O2 )= 0,045*32=1,44 g.


answer: 1,44 g.